OpenStudy (anonymous):

PLEASE HELP WILL GIVE A MEDAL Please help and explain all steps! Factor completely: a^2 - a - 20

OpenStudy (anonymous):

The 9th space has the number 4!

OpenStudy (p0sitr0n):

a^2-a-20=a^2-5a+4a+(-5)(4)=a(a+4)-5(a+4)=(a-5)(a+4)